OpenStudy (anonymous):

If 2x+3y=a and -x+5y=6b, express x and y in terms of a and b. (Vectors)

OpenStudy (anonymous):

@asnaseer hi, do you know how to solve this question? :$

OpenStudy (anonymous):

Somehow I have to use substitution? i'm not so sure.. :|

OpenStudy (asnaseer):

use the second equation to express x in terms of y and b

OpenStudy (asnaseer):

then substitute that into the first equation and rearrange to get y in terms of a and b

OpenStudy (asnaseer):

so lets start with the first step - can you rearrange the second equation to get an expression for x in terms of y and b?

OpenStudy (anonymous):

so sorry for the late reply but is it: x=5y-6b?

OpenStudy (asnaseer):

yes, now use this to replace x in the first equation

OpenStudy (anonymous):

so, 2(5y=6b)...

OpenStudy (anonymous):

*-6b

OpenStudy (asnaseer):

yes, except that should be 2(5y-6b)... use that to find an expression for y in terms of a and b

OpenStudy (anonymous):

what do you mean?

OpenStudy (asnaseer):

ok, first substitute the expression for x into the first equation - what do you get - please write out the entire expression.

OpenStudy (anonymous):

2(5y-6b)+3y=a

OpenStudy (asnaseer):

good - now expand the brackets

OpenStudy (anonymous):

10y-12b+3y=a

OpenStudy (asnaseer):

great! - now gather the terms involving y and move everything else to the right-hand-side of the equals sign

OpenStudy (anonymous):

so, 13y=a+12b?

OpenStudy (asnaseer):

yes, now divide both sides by 13 to get an expression for y in terms of a and b

OpenStudy (anonymous):

and that's it? o.O

OpenStudy (asnaseer):

yes - that gives you y in terms of a and b. then substitute that expression for y back into the other equation you found: x=5y-6b to find x in terms of a and b as well.

OpenStudy (anonymous):

ohh, thank you! btw, how did you know that we had to use substitution (aside from telling you)? o.o

OpenStudy (asnaseer):

it was from "you telling me" :)

OpenStudy (asnaseer):

is that how the question was originally worded?

OpenStudy (anonymous):

ohh.. what if I didn't tell you-- how would you have known? o.O

OpenStudy (asnaseer):

it would have been fairly obvious. it is just plain algebra.

OpenStudy (anonymous):

aww man, okay thank you!

OpenStudy (asnaseer):

yw :)